solving equations that contain fractions. 6x-1 divided by 5 minus 2x divided by 3 = 3
I used the LCM which is 15, then I used the distrubutive property for both fractions. From then on I don't know what to do, if I am doing it right at all, I don't know. Please help.

Answer by stanbon(75887)

solving equations that contain fractions. 6x-1 divided by 5 minus 2x divided by 3 = 3
(6x-1)/5 -(2x)/3 = =3
Multiply through by 15 to get:
3(6x-1)-5(2x) = 45
18x-3-10x=45
8x=48
x=6
Cheers,
Stan H.